OpenStudy (anonymous):

What are the vertex and x-intercepts of the graph of the function below? y =(x + 4)(x + 2) A. Vertex: (1, -5); Intercepts: x = -4, 2 B. Vertex: (1, -9); Intercepts: x = 4, -2 C. Vertex: (-4, 2); Intercepts: x = 3, -3 D. Vertex: (-3, -1); Intercepts: x = -4, -2

OpenStudy (triciaal):

x-intercept means when y = 0 when y = 0 (x +4) = 0 or (x +2) = 0 find your x values

OpenStudy (anonymous):

Ah, I'm sorry for such a late reply. I really don't know these at all. :L

OpenStudy (triciaal):

option 1 x^2 + 6x + 8 = 0 a=1, b = 6, c = 8 x-value of vertex = -b/2a y-value = f(-b/2a)

OpenStudy (anonymous):

Is the x-intercept by chance -3?

OpenStudy (triciaal):

option 2 how would you complete the square?

OpenStudy (triciaal):

no the vertex at x = -3

OpenStudy (triciaal):

the x-intercept when y = 0

OpenStudy (triciaal):

read the very first response

OpenStudy (anonymous):

My apologies, I'm very bad at this. Would it be option D?

OpenStudy (triciaal):

yes

OpenStudy (anonymous):

Thank you :3

OpenStudy (triciaal):

please review the process welcome

OpenStudy (anonymous):

I'll definitely be sure to keep this in mind, thank you! ❤

OpenStudy (triciaal):

no problem